The Hartnett Corporation manufactures baseball bats with Pudge Rodriguez’s autograph stamped on them. Each bat sells for $59 and has a variable cost of $31. There are $42,840 in fixed costs involved in the production process.

a. Compute the break-even point in units.
  



b. Find the sales (in units) needed to earn a profit of $15,960.
Answer #1

Contribution margin=Sales-Variable cost

=(59-31)=$28 per unit

a.Breakeven point=Fixed costs/Contribution margin

=42,840/28

=1530 units

b.Target Contribution margin=Fixed cost+Target profits

=(42,840+15,960)=$58800

Hence target sales=58800/28

=2100 units.